A uniform electric field E = 8000 N/C passing through a flat square area A = 10 m2. Determine the electric flux.

1
Expert's answer
2022-02-28T10:42:50-0500

The electric flux

"\phi=EA\cos\theta\\
=\rm 8000\: N/C*10\: m^2*\cos 0^\circ=8*10^4\: N\cdot m^2/C"